OpenStudy (anonymous):

Help:Convert the fraction to a decimal 4/25 A. 0.16 B. 0.425 C. 1.6 D. 4.25

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@viop9 thanks

OpenStudy (anonymous):

@kia55 np

OpenStudy (anonymous):

\[\frac{ 4 }{ 25 }=\frac{ x }{ 100 }=>25x=400=>x=16=>\frac{ 16 }{ 100 }=0.16\]

OpenStudy (anonymous):

@hwtrain Thanks :)